### v1

# A Time Series Model for Three Asset Classes used in Financial Simulator

Abstract.
We create a dynamic stochastic general equilibrium model for annual returns of three asset classes: the USA Standard & Poor (S&P) stock index, the international stock index, and the USA Bank of America investment-grade corporate bond index. Using this, we made an online financial app simulating wealth process. This includes options for regular withdrawals and contributions. Four factors are: S&P volatility and earnings, corporate BAA rate, and long-short Treasury bond spread. Our valuation measure is an improvement of Shiller’s cyclically adjusted price-earnings ratio. We use classic linear regression models, and make residuals white noise by dividing by annual volatility. We use multivariate kernel density estimation for residuals. We state and prove long-term stability results.
1. Introduction
1.1. Index levels and earnings
A classic way to find whether the stock market is priced fairly is to compare index levels with earnings. The classic cyclically adjusted price-earnings (CAPE) ratio, introduced by Robert Shiller, measures how reasonably the stock market (measured by the Standard & Poor 500 and its predecessors) is priced relative to fundamentals (earnings). It is computed by dividing the current index level by annual earnings, averaged over the previous 10 years. Thie latter, called cyclical averaging, is done to smooth out wild short-term fluctuations in earnings, and take the data from both recessions and expansions. Historically, this ratio is between 5 and 45. Its long-term median value is 17. If it is much larger than that, then the market is overpriced and is likely to crash soon, or at least not grow as fast as before. See the classic article [6] or the monograph [16], or a more recent article [12] and a textbook [10, Chapter 8, Sections 4, 6].
Robert Shiller did a remarkable job compiling historical data from 1871 for Standard & Poor 500 index and its predecessors. Three main examples of such overprived markets in the USA history are: late 1920s followed by the Great Depression; the boom of 1960s followed by the stagflation and energy crises of 1970s; and the dotcom bubble in the late 1990s, followed by another stock crash and a lost decade of the 2000s.
However, a surprising observation was made based on recent data: In the late 2010s and now (2025 as of this writing), the CAPE ratio is very high compared to historical average: Now, it is around 40. In fact, it is close to its all-time peak in the late 1990s at the height of the dotcom bubble. However, the market crash failed to materialize so far. There was a short-timed crash in early 2020 during the lockdowns, but the market has recovered quickly. Did the CAPE lose its predictive value?
1.2. The new valuation measure
In our previous article [13], we modified the CAPE. We think that CAPE lost its importance because dividend payout decreased. Historically, about 70% of earnings of stocks were paid as dividends to shareholders. However, in the last few decades, this decreased to 40% on average. Instead of paying earnings as dividends, executives prefer to use earnings for stock buybacks. This increases the stock value.
Reasons are still disputed in the academic community, but likely include tax issues, [10, Chapter 8, Section 4], [1, Chapter 12, Section 5], as well as incentives for executives who get paid in bonuses of stocks in their own company, [1, Chapter 15]. During stock buybacks, overall earnings remain the same, but index levels increases. This, in turn, boosts the CAPE ration, but does not make the market fundamentally overpriced.
To capture this, in [13] we compared annual total returns with annual earnings growth. In our research, annual returns are geometric (not arithmetic, which are commonly quoted); total (including dividends); and nominal (not adjusted for inflation). Annual earnings growth is measured as logarithmic change, and also nominal (without inflation adjustments). The long-term difference is 4–5% per year. Cumulative total returns minus earnings have linear trend with slope 4–5%. After detrending, this becomes mean-reverting: Modeled as a classic autoregression of order 1. This detrended quantity is our new valuation measure. Our previous research showed significant predictive power for next year’s total returns.
(a) CAPE vs New Measure
(b) Volatility
Figure 1. On the left: Shiller cyclically-adjusted price-earnings (CAPE) ratio on the log scale versus the new valuation measure proposed in [13]. They track each other before 2000, but after that they diverge. On the right: annual realized volatility for the Standard & Poor. This graph is taken from [13].
This new valuation measure more or less tracks the CAPE in the 19th and 20th centuries. The new valuation measure reproduces these 3 peaks in the 1920s, 1960s, and 1990s mentioned above. However, after the dotcom bubble crashed in the 21st century, these two measures diverge. The CAPE shoots upward during the 2010s, but the new measure stays close to its historical average. This is a reason to believe that the stock market now, in fact, is reasonably priced, and there is no bubble. Comparison of the Shiller CAPE and the new valuation measure is in Figure 1 (A).
1.3. Market volatility
Assume we model earnings growth separately. Then we can model earnings and the index together, using this valuation measure. Indeed, this measure is the difference between total returns and earnings growth. Knowledge of earnings growth gives us knowledge of total returns. But it is much harder to model earnings growth terms: These are not independent identically distributed (IID), that is, not a (strong) white noise, judging by the autocorrelation function (ACF). In fact, the returns $Q(t)$ themselves have ACF close to zero, but the squared returns $Q^{2}(t)$ or their absolute values $|Q(t)|$ do not have zero ACF. For background, see the classic references [3, Chapter 7], especially Figures 7.1–7.3. A solution comes from our previous article [14], using the volatility of the S&P index. For background on such stochastic volatility models, see [3, Chapter 7, Section 4] or [9, Chapter 3, Section 1] (especially Figures 3.1, 3.2).
We used monthly average volatility index (VIX), computed daily by the Chicago Board of Options Exchange (CBOE) from implied volatility (computed using prices of European options traded on this exchange). See background in [1, Chapter 7, Section 2, Figure 7.3]; also [1, Chapter 10, Section 1] for volatility-driven strategies; or [10, Chapter 8, Section 4, Figure 8.5]. Dividing monthly S&P 500 returns by this VIX made these returns IID Gaussian. Unfortunately, applying this methodology directly to our setting is problematic: We have annual (not monthly) data, and the VIX data goes back only as far as 1990 (or 1986 for the version of VIX for S&P 100, a close relative of S&P 500).
To resolve this difficulty, we switched to annual realized volatility (as opposed to implied volatility): the standard deviation of daily index fluctuations (log changes) for S&P 500 and its predecessor S&P 90, within each year 1928–2024. This range of years is more limited than 1871–2024 from Robert Shiller’s data library. Still, it includes the three great bubbles. See the graph on Figure 1 (B).
What is more, we improved data for S&P 500 itself: We take end-of-year (December 31 or the last working day of each year) index level (available from 1927) instead of the average January daily close data (available from Robert Shiller’s data library from 1871). This is important, since monthly average is not directly investable, but the end-of-year level is.
Annual returns are not IID, but dividing them by annual volatility makes them IID. Same for the annual earnings growth. The volatility itself can be modeled as an autoregression of order 1, if you take the logarithm of volatility. Thus we described two factors influencing the stock market returns: Volatility and the new valuation measure (computed using earnings).
1.4. Bond rates and spreads
However, there exist other factors influencing the S&P 500 total returns. Bond rates are very important. Bond and stock markets compete for the capital of investors. An increase in bond rates makes the bond market more attractive, thus reducing demand for stocks. As a benchmark, we took the BAA rate for US corporate bonds. This Moody’s rating is less than the perfect AAA, and is on the lower range for investment-grade bonds. Still, these bonds are relatively safe from defaults. We take average daily rate in December of 1927–2024. See the classic article [2] for reference. But here we take the BAA rate instead of the short-term Treasury rate, since it is easier to model it as an autoregression. Also, see textbooks [1, Chapter 9, Section 4] and [10, Chapter 9, Section 4].
(a) Spread
(b) BAA rate
Figure 2. The long-short (10 year minus 3 month) Treasury spread, and the Moody’s BAA rate: Average daily December data.
Yet another factor is the long-short Treasury spread, sometimes called term spread. Usually, long-term Treasury bonds have higher rate than short-term Treasury bills. This is natural, since committing money for a longer time is more risky and demands compensation in the form of a higher rate. Occasionally, however, this comparison inverts: Long-term Treasuries have lower rate than short-term Treasuries. This phenomenon is called the inverted yield curve. This usually happens when the market anticipates an impeding recession.
The Federal Reserve System has direct influence over short-term rates via its standard open-market operations (however, recently it sometimes intervenes in long-term bond market as well). When a recession strikes, the Federal Reserve lowers short-term rates. However, the long-term rate now combines the current short-term rate and the anticipated future short-term rates. Indeed, an alternative to buying long-term bonds is reinvesting many times in short-term bonds. In sum, when long-term rates are lower than the short-term rates, then a recession and a stock market crash are likely to happen soon. For the term spread, we consider December average daily data for 10 years minus 3 month Treasury rates. For reference, see for example [8], and more recent textbooks [1, Chapter 9, Section 3], especially Figure 9.9; or [10, Chapter 9].
We model both BAA rate and the term spread as autoregression of order 1. For the BAA rate, we take its logarithm to prevent it from being negative, similarly to the volatility. See the graphs in Figure 2.
1.5. Our main contributions
In our previous research, we proved long-term stability only for the valuation measure. Here, we have a feedback loop: The stock market returns are modeled using linear regression with factors including the new valuation measure, but the latter in return is computed using earnings growth and the stock market returns. Thus we need to state and prove a non-trivial long-term stationarity result. It does not follow immediately from our assumptions.
Long-term stability is a desirable property for any financial econometric model. We create a time series model which is called DSGE: dynamic (time series equations), stochastic (each time step includes random white noise, which is IID), and general equilibrium (stationary). See, for example, a textbook [7].
Also, the simulator includes the international stocks and USA bonds. This is why we model this using linear regression as well, using rates and volatility. We provide full statistical analysis and prove the stationarity for this extended model. We apply this simulator for retirement modeling, to test the classic 4% withdrawal rule.
The Excel data and Pyhon code for the fitting of the model, analysis of residuals, and the web app itself, is available in the repository asarantsev/sim Web app has back-end written in Python, using the Flask framework, with bare HTML for the front-end. The web address of the simulator itself is asarantsev.pythonanywhere.com
1.6. Organization of the article
In Section 2, we describe the raw data, and the preliminary processing. We discuss the sources and forats of data in detail. Section 3 is devoted to methodology of white noise analysis. In Section 4, we describe the simple model with only two factors: volatility and BAA rate. In Section 5, we describe in detail the new valuation measure, and use it for combined model of S&P returns and earnings. In Section 6, we consider the complete model. For each model, we state and prove the main stability result. Section 7 describes the online financial simulator based on the model from Section 6.
2. Data Description
2.1. Preliminary remarks
For this background, see [9, Chapter 1, Section 1]. We consider total returns (including dividends or coupon payments, not just price changes), nominal (not inflation-adjusted) rather than real (inflation-adjusted), and geometric (log returns) rather than arithmetic: If 1\$ turned into 2\$ the returns are $\ln 2$ not 100%. Arithmetic and geometric returns $r_{a},r_{g}$ are related as $r_{g}=\ln(1+r_{a})$. However, if $r_{a},r_{g}$ are small, then $r_{a}\approx r_{g}$.
Using geometric returns allows us to aviod compound interest: If $r_{1}$ and $r_{2}$ are geometric returns of this asset class in years 1 and 2, then $r_{1}+r_{2}$ are geometric returns of this asset class in two years. However, arithmetic returns are better for computing portfolio returns: If $r_{1}$ and $r_{2}$ are arithmetic returns of asset classes 1 and 2 in year $t$, and if $\pi_{1}$ and $\pi_{2}$ are portfolio weights for a constant-weighted portfolio, then $\pi_{1}r_{1}+\pi_{2}r_{2}$ are arithmetic returns of this portfolio in year $t$.
2.2. Online data resources
We used the following web sites:
  • Robert Shiller’s Yale University financial data library (Shiller): shillerdata.com
  • Federal Reserve Economic Data (FRED): fred.stlouisfed.org
  • Yahoo Finance: finance.yahoo.com
  • Novel Investor web site: novelinvestor.com/historical-returns
2.3. Raw data
All data is nominal.
  • Daily close December 30, 1927–December 31, 2024 for Standard & Poor (S&P): Standard & Poor 500, 1957–2024, and its predecessor, Standard & Poor 90, 1927–1956. Source: Yahoo Finance ˆSPX series
  • Annual S&P earnings $E(t)$, 1918–2024, and dividends, 1928–2024. Source: Shiller, file ie_data.xls columns named Dividend D and Earnings E.
  • Moody’s BAA corporate bond rate $R(t)$, average daily close December, 1927–2024. Source: FRED, BAA series
  • Treasury long rates, average daily close December, 1927–2024. Source: FRED series DGS10, 10-year Treasury rate
  • Treasury short rates, average daily close December, 1927–2024. Source: FRED series TB3MS, 3-month Treasury rate, 1934–2024, FRED series M1329AUSM193NNBR, 3- and 6-month short-term certificates of deposit, 1927–2024
  • Bank of America (BofA ICE) Corporate Bonds Total Return Index Value, $\mathcal{U}(t)$ end-of-year for 1972–2024. Source: FRED series BAMLCC0A0CMTRIV
  • International annual total returns: Morgan Stanley Capital International (MSCI) Europe, Australasia & Far East (EAFE), 1970–2024, index; MSCI Emerging Markets (EM), 1988–2024, index. Source: Novel Investor
Using this raw data, we compute the following:
  • From the daily S&P index, we compute daily log changes of S&P. Within each year 1928–2024, we compute the standard deviation of these daily log changes. This is the annual S&P (realized) volatility $V(t)$, 1928–2024.
  • We subtract Treasury short rates from Treasury long rates, and get Treasury term spread $S(t)$, or long-short spread, 1927–2024.
  • Total S&P annual returns $Q(t)$ are computed from end-of-year S&P index and its annual dividends, 1928–2024.
  • For the averaging window $L=1,\ldots,10$, trailing averaged earnings $\overline{E}(t)$ of year $t$ are the average of $E(t),\ldots,E(t-L+1)$ annual earnings, for 1927–2024.
  • We compute earnings growth $G(t)$ as log changes in annual earnings, 1928–2024, and averaged earnings growth $\overline{G}(t)$ as log changes in annual averaged earnings, 1928–2024.
  • We compute international stock returns $I(t)$ as EAFE from 1970–1987 and the 60/40 portfolio of EAFE/EM from 1988–2024, thus having the data 1970–2024. The 60/40 choice is arbitrary and can be changed in further research.
  • Compute corporate bond returns $B(t)$ as log changes in their Index Value, 1928–2024.
We also provide the data on end-of-year annual S&P index, earnings, dividends, bond rates, and $I(t),\mathcal{U}(t)$ from above, on our web site asarantsev.github.io and in GitHub repository github.com/asarantsev/sim
2.4. Asset classes and return factors
Let us describe in detail the three asset classes. We denote their total returns by letters.
  • USA stocks $Q$, measured by the S&P index, 1928–2024.
  • International stocks $I$, measured by the MSCI indices: The EAFE index, 1970–1987, and the 60/40 portfolio of the EAFE/EM indices, 1988–2024.
  • USA investment-grade corporate bonds $B$, measured by BofA ICE Corporate Bond Index, 1973–2024.
The four factors which we use for modeling in this article:
  • Annual S&P volatility $V$
  • Moody’s BAA corporate bond rate $R$
  • Long-short term spread $S$
  • Annual S&P earnings $E$
3. White Noise Analysis Methodology
The code is Monte-Carlo.py and results are in Table 1. We shall use Shapiro-Wilk and Jarque-Bera normality tests, as well as skewness, kurtosis, and the ACF. Using Monte Carlo simulation, we provide critical values for IID $x_{1},x_{2},\ldots,x_{N}\sim\mathcal{N}(\mu,\sigma^{2})$ for the following statistics: empirical skewness, empirical kurtosis (with both theoretical skewness and kurtosis equal zero for normal distribution; and the $L_{1}$ norm of the empirical ACF $\hat{\rho}(k)$ with lag $k$. Here, $\overline{x}$ and $s$ are empirical mean and standard deviation of the data $x_{1},\ldots,x_{N}$.
$L_{1}=\sum_{k=1}^{5}|\hat{\rho}(k)|,\quad\hat{\rho}(k)=\frac{1}{Ns^{2}}\sum\limits_{j=1}^{N-k}(x_{j}-\overline{x})(x_{j+k}-\overline{x}).$
For background on white noise and normality tests, see [3, Chapter 1, Section 6]. Also, for white noise testing, see [9, Chapter 1, Sections 2, 4], and for normality testing, see [9, Chapter 1, Section 5]; finally, see [9, Chapter 2, Section 6], and the result (2.10) there.
Size NN $p$ Skew Kurt L1
100 95% 0.47 0.76 0.63
100 99% 0.63 1.38 0.76
50 95% 0.64 0.98 0.88
50 99% 0.88 1.86 1.08
Table 1. Critical values for skewness, kurtosis, and $L_{1}$ norm of the first 5 lags of the ACF for original and absolute values of standard normal variables.
Note that we perform the ACF computation both for $x_{1},\ldots,x_{N}$ and for $|x_{1}|,\ldots,|x_{N}|$. The necessity of doing this for $x_{1}^{2},\ldots,x_{N}^{2}$ and not just for $x_{1},\ldots,x_{N}$ is shown in [9, Chapter 3, Section 1] (especially Figures 3.1, 3.2), mentioned in the Introduction. To avoid computing higher moments, which might lead to increased standard error, we use absolute values instead of squares: $|x_{1}|,\ldots,|x_{N}|$. But both $x_{1},\ldots,x_{N}$ (the original values) and $|x_{1}|,\ldots,|x_{N}|$ (absolute values) give us the same critical values of the $L_{1}$ statistic.
We recall the standard time series results: Under the white noise hypothesis, we have the following weak convergence: $(\hat{\rho}(1),\ldots,\hat{\rho}(k))\Rightarrow\mathcal{Q}_{k}$, where $N\to\infty$, where $\mathcal{Q}_{k}$ is the standard $k$-variate normal distribution. From here, it is easy to derive asymptotic results for $L_{1}$ and $L_{2}$. But we decided to make Monte Carlo simulations and compute critical values by hand, since we are not sure about the rate of convergence. After all, this was in part what motivated the replacement of the standard $L_{2}$ Box-Pierce test with the Ljung-Box test, see the classic article [11].
4. The Simplified Model
In this section, we consider the model of returns of three asset classes, but using only two of the four above factors: S&P volatility $V$ for stocks and rate $R$ for bonds. This was similar to the model implemented in previous versions of the simulator. We discuss this model in detail to illustrate our main ideas. The Python code is simple.py
4.1. Stock returns
It is reasonable to model domestic stock returns $Q$ by terms which are IID but not Gaussian. But after normalizing (dividing by volatility), $Q/V$ can be modeled as IID Gaussian, see Figure 3. See also Table 2.
Somewhat suprpisingly, we see it is reasonable to model international returns $I$ as IID Gaussian even without normalizing. But we think it is a statistical illusion because we do not have that much annual data for international stocks. We still decide to normalize international returns and consider $I/V$.
It is remarkable that we can normalize international stock returns by dividing by domestic volatility. This works for international as well as for domestic stocks. We explain it in the following way: World equity markets are interconnected, and when volatility of domestic markets increases, the same happens with international markets.
returns stdev skew kurt SW JB L1O L1A
$Q$ 0.1853 -0.971 1.082 0.000 0.000 0.327 0.265
$Q/V$ 0.0197 0.228 -0.211 0.115 0.601 0.558 0.497
II 0.2120 -0.617 1.015 0.168 0.054 0.488 0.563
$I/Q$ 0.0236 0.604 -0.129 0.049 0.184 0.392 0.207
Table 2. Descriptive statistics of domestic and international stock returns before and after normalizing. Skew = skewness, Kurt = kurtosis, SW and JB = Shapiro-Wilk and Jarque-Bera $p$-values, L1O and L1A = L1 norm for the first 5 lags of the ACF for $Z$ and $|Z|$
(a) $Q$
(b) $Q$
(c) $|Q|$
(d) $Q/V$
(e) $Q/V$
(f) $|Q/V|$
Figure 3. Top: The quantile-quantile plot for $Q$, and the ACF for $Q$ and for $|Q|$, where $Q$ is S&P returns. This shows $Q$ are IID but not Gaussian. Bottom: The quantile-quantile plot for $Q/V$, and the ACF for $Q/V$ and for $|Q/V|$, where $Q$ is total S&P returns. This shows $Q/V$ are IID Gaussian.
4.2. Stock market volatility
Following [14] with monthly implied volatility analysis, annual volatility on the log scale is modeled as an autoregression of order 1:
(1) $\ln V(t)=\alpha_{V}+\beta_{V}\ln V(t-1)+Z_{V}(t).$
We use the log scale to make sure the volatility is always positive. Fitting (1), we get: $\alpha_{V}=0.847850$ and $\beta_{V}=0.620146$. We reject the hypothesis $\beta_{V}=1$ (the random walk), with $p<0.1\%$. Analysis of residuals shows that they are IID but not Gaussian, see Table 3.
4.3. Bond rates and returns
Returns $B(t)$ of bonds are modeled using bond rates $R(t)$. The simplest approximation of bond returns during year $t$ is
(2) $B(t)\approx 0.01\cdot R(t-1).$
This is true if we had arithmetic returns $B(t)$; but geometric returns are close to arithmetic returns if both are small. A bond with 5% rate at end of year t1t-1 will return approximately 5% in year $t$. However, all bonds have interest rate risk: Changes in rates during year $t$ affect the price of this bond. Sensitivity to this change is called duration. An increase in rates leads to a decrease in prices. This duration is approximately the average of dates for payments (coupon payments, usually every 6 months; and the principal at maturity), weighted by the size of these payments. For example, a zero-coupon bond has duration equal to its maturity. But corporate bonds usually do have coupons. Assuming $d_{B}$ is this duration, we add the following term to (2):
(3) $B(t)\approx 0.01\cdot R(t-1)-d_{B}(R(t)-R(t-1)).$
Finally, corporate bonds do have a small risk of default: Not paying in full their coupons or the principal. This stands in contrast with Treasury bonds, which have interest rate risk but not default risk. We need to substract a constant from the right-hand side of (3):
(4) $B(t)-0.01\cdot R(t-1)=-a_{B}-d_{B}(R(t)-R(t-1))+Z_{B}(t).$
Results are: $a_{B}=0.016611$ and $d_{B}=0.055884$. The parameter $m_{0}$ stands for the duration: sensitivity of corporate bond prices to changes in interest rates. Here, the duration is about $5.5$ years. The parameter $a_{B}$ measures default rates. The $p$-values for the Student $T$-test are less than $0.1\%$.
Finally, the BAA corporate bond rate is also modeled as an autoregression of order 1, after taking the logarithm to ensure they stay positive:
(5) $\ln R(t)=\alpha_{R}+\beta_{R}\ln R(t-1)+Z_{R}(t).$
Fitting (5), we get: $\alpha_{R}=0.643694$ and $\beta_{R}=0.539518$. Note that the $p$-value for $\beta_{R}=1$ using Student $T$-test is $p=9.4\%$. Thus we fail to reject the null hypothesis that $\beta_{R}=1$, that is, the process $\ln R$ is a random walk.
4.4. Combined simple model
We combine equations  (1), (5), (4) with the observations about normalized returns:
[…]
4.5. Including duration for stocks
Bond returns are influenced by the duration, discussed earlier in (4). We can also consider stock duration: Regress $Q(t)$ and $I(t)$ versus R(t)R(t1)R(t)-R(t-1). However, we do not wish to write a simple linear regression:
(7) $Q(t)=c_{Q}-d_{Q}(R(t)-R(t-1))+Z_{Q}(t),$
since we normalized $Q$ by dividing it by volatility. Only this makes data Gaussian. So we rewrite (7) Instead of (6), we get:
(8) $\frac{Q(t)}{V(t)}=\frac{c_{Q}}{V(t)}-d_{Q}\frac{R(t)-R(t-1)}{V(t)}+Z_{Q}(t).$
But usually, a linear regression has an intercept. We add such intercept to (8) and get:
(9) $\frac{Q(t)}{V(t)}=a_{Q}-d_{Q}\frac{R(t)-R(t-1)}{V(t)}+\frac{c_{Q}}{V(t)}+Z_{Q}(t).$
We can rewrite (9) as $Q(t)=a_{Q}V(t)-d_{Q}(R(t)-R(t-1))+c_{Q}+V(t)Z_{Q}(t)$. The same can be done for returns of international instead of domestic stocks. Combining all these, we get:
[…]
We have coefficients
$a_{Q}$ $=0.2121,\quad d_{Q}=0.0621,\quad c_{Q}=-0.0109,$
$a_{I}$ $=0.2829,\quad d_{I}=0.0508,\quad c_{I}=-0.0189.$
The coefficients $d_{Q},c_{Q},d_{I},c_{I}$ have Student $T$-test $p$-values $0.1\%,0.8\%,2\%,2\%$. Analysis of residuals show that all five series of residuals are IID, but only $Z_{Q},Z_{I},Z_{B}$ are Gaussian.
Reg Size Stdev Skew Kurt SW JB L1O L1A
$Z_{V}$ 96 0.3644 0.590 0.057 0.009 0.061 0.401 0.237
$Z_{R}$ 97 0.1357 0.814 1.409 0.007 0.000 0.177 0.360
$Z_{Q}$ 97 0.0149 0.226 0.068 0.420 0.655 0.480 0.435
$Z_{I}$ 55 0.0183 -0.059 -0.501 0.754 0.738 0.401 0.519
$Z_{B}$ 52 0.0263 0.193 0.238 0.857 0.800 0.878 0.706
Table 3. Analysis of regression residuals. Skew = skewness, Kurt = kurtosis, SW and JB = Shapiro-Wilk and Jarque-Bera $p$-values, L1O and L1A = L1 norm for the first 5 lags of the ACF for $Z$ and $|Z|$, Length = the size of $Z$
The empirical correlation matrix of $(Z_{V},Z_{R},Z_{Q},Z_{I},Z_{B})$ is
$\Sigma_{5}:=\begin{bmatrix}1&-0.3&-0.12&-0.1&0.07\\ 0.3&1&-0.14&-0.04&-0.15\\ -0.12&-0.14&1&0.29&0.41\\ -0.1&-0.04&0.29&1&-0.07\\ 0.07&-0.15&0.41&-0.07&1\\ \end{bmatrix}$
Theorem 1.
If $\beta_{V},\beta_{R}\in(0,1)$, and if $(Z_{Q}(t),Z_{I}(t),Z_{B}(t),Z_{V}(t),Z_{R}(t))$ are IID with mean zero, the system (10) has a unique stationary version.
5. Modeling Earnings Growth and the New Valuation Measure
5.1. The valuation measure
In this subsection, we mimic the econometric analysis of [13, Section 2.1]. Define the wealth process as $W(t)=\exp(Q(1)+\ldots+Q(t))$, with initial wealth $W(0)=1$. We use $L$-year averaged trailing earnings $\overline{E}$ for $L=10$, and their growth:
(11) $\overline{E}(t)=\frac{1}{L}\sum\limits_{s=0}^{L-1}E(t-s),\quad\overline{G}=\ln\overline{E}(t)-\ln\overline{E}(t-1).$
We compare wealth with averaged trailing earnings:
(12) $\ln W(t)-\ln\overline{E}(t)=Q(1)+\ldots+Q(t)-\overline{G}(1)-\ldots-\overline{G}(t)-\ln E(0).$
As discussed in the Introduction, long-term growth of $\ln W$ is 6–7% and long-term growth of $\ln E$ (and therefore $\ln\overline{E}$) is 2%. Therefore, the long-term growth $c$ of such difference from (12) is 4–5%. After detrending (subtracting the linear trend $ct$, which is to be determined), we get the new valuation measure
(13) $H(t)=\ln W(t)-\ln\overline{E}(t)-ct=\sum\limits_{s=1}^{t}(Q(s)-\overline{G}(s))-ct.$
We model this quantity $H(t)$ from (13) using autoregression of order 1:
(14) $H(t)-h=b(H(t-1)-h)+U(t).$
Here, $h$ plays the role of the long-term average of this new valuation measure $H$, and $b$ is the mean-reverting slope coefficient of this autoregression.
Figure 4. The new valuation measure, computed in this article, using end-of-year S&P data instead of January average S&P data, and 10-year instead of 5-year averaging window. It closely tracks but is slightly different from the new valuation measure on Figure 1. However, it shows the three main peaks in 1920s, 1960s, and 1990s. As of 2024, the stock market is not overvalued.
Using (13), we rewrite (14) below to fit as an ordinary linear regression.
(15) $Q(t)-\overline{G}(t)=\alpha+\beta(t-1)-\gamma\sum\limits_{s=1}^{t-1}(Q(s)-\overline{G}(s))+U(t).$
coefficient estimate Student $p$-value stderr
$\alpha$ 0.023893 51.6% 0.037
$\beta$ 0.008608 0.3% 0.003
$\gamma$ 0.190133 0.2% 0.060
Table 4. Regression results for (15). Coefficients $\beta$ and $\gamma$ are statistically significant. Innovations $U$ are IID but not Gaussian.
Results of fit of (15) are shown in Table 4. Solving for $h,b,c$ in (13), (14), similarly to [13, (6)], we get:
(16) $b=1-\gamma=0.819867,\quad c=\frac{\beta}{\gamma}=0.04527,\quad h=\frac{\alpha-c}{\gamma}=-0.1124.$
We reject the hypothesis $b=1$, so the autoregression (14) is mean-reverting, not a random walk. Residuals $U$ are IID but not Gaussian, judging by statistics in Table 5.
5.2. Earnings growth
For annual earnings $E(t)$, their growth $G(t)=\ln E(t)-\ln E(t-1)$ can be also modeled using annual volatility. Terms $G(t)$ are not IID. But dividing these growth terms by annual market volatility $V$ makes them white noise: $G(t)/V(t)$ are IID and Gaussian, see Figure 5. To us, it is surprising that dividing $G$ by $V$ makes these IID. After all, $V$ is the volatility of daily index changes, not annual earnings. However, this method does work. The quantity $V$ plays universal role for stock markets: It improves (makes closer to IID) S&P returns, international returns, and S&P earnings.
(a) $G$
(b) $G$
(c) $|G|$
(d) $G/V$
(e) $G/V$
(f) $|G/V|$
Figure 5. Top: The quantile-quantile plot for $G$, and the ACF for $G$ and for $|G|$, where $G$ is S&P returns. This shows $G$ are not IID and not Gaussian. Bottom: The quantile-quantile plot for $G/V$, and the ACF for $G/V$ and for $|G/V|$, where $G$ is total S&P returns. This shows $G/V$ are IID Gaussian.
5.3. Combined model of S&P returns and earnings
This gives us the following model for S&P returns and earnings. This is an improvement of our previous article [13] where we did not model earnings growth. Therefore, our model in that article was not complete. Combine (12), (13), (14) with definitions of $G$ and $\overline{G}$.
Reg Size Stdev Skew Kurt SW JB L1O L1A
$Z_{V}$ 96 0.3644 0.590 0.057 0.009 0.061 0.401 0.237
$Z_{G}$ 97 0.0218 0.614 2.903 0.000 0.000 0.474 0.253
$U$ 97 0.1767 -0.816 1.102 0.003 0.000 0.291 0.608
Table 5. Analysis of regression residuals. Skew = skewness, Kurt = kurtosis, SW and JB = Shapiro-Wilk and Jarque-Bera $p$-values, L1O and L1A = L1 norm for the first 5 lags of the ACF for $Z$ and $|Z|$, Length = the size of $Z$.
We consider modeling $G/V$ as IID. But we also add the intercept, similarly to switching from (8) to (9). We get:
(17) $\frac{G(t)}{V(t)}=g+\frac{c_{G}}{V(t)}+Z_{G}(t)\Leftrightarrow G(t)=gV(t)+c_{G}+V(t)Z_{G}(t).$
Combning (11), (13), and (14), we get the model:
[…]
We have $c_{G}=0.123275,g=-0.005788$. The $p$-values for the Student $T$-test are $0.8\%$ and $32\%$, respectively. The correlation matrix of residuals is
$\Sigma_{3}=\begin{bmatrix}1&-0.18&-0.42\\ -0.18&1&0.1\\ -0.42&0.1&1\\ \end{bmatrix}$
All these computations in this section are done in Python code bubble.py
Theorem 2.
If $\beta_{V},b\in(0,1)$ and $(Z_{G}(t),Z_{V}(t),U(t))$ are IID with mean zero, then the system (18) has a stationary version.
Proof.
Stationarity of $V$ and $H$ is immediate. Clearly, $G$ is also stationary. Finally, let us show stationarity of $Q$. We can rewrite (13) and (14) as
$Q(t)$ $=H(t)-H(t-1)+c+\overline{G}(t);$
$\overline{G}(t)$ $=\ln\overline{E}(t)-\ln\overline{E}(t-1).$
It suffices to prove $\overline{G}(t)$ is stationary. But we can rewrite
$\overline{G}(t)$ $=\ln\frac{\overline{E}(t)}{E(t-L)}-\ln\frac{\overline{E}(t-1)}{E(t-L)}=\ln\sum\limits_{s=0}^{L-1}\frac{E(t-s)}{E(t-L)}-\ln\sum\limits_{s=0}^{L-1}\frac{E(t-s-1)}{E(t-L)}.$
The terms E(ts)/E(tL)E(t-s)/E(t-L) are stationary for each $s$, and are jointly stationary for $s=0,\ldots,L$, for the simple reason that these all are products of $G(t),\ldots,G(t-L+1)$. For example, $E(t)/E(t-L)=G(t)\cdot\ldots\cdot G(t-L+1)$. This proves stationarity for $\overline{G}(t)$, and therefore for $Q(t)$. ∎
5.4. Lag selection
Try other averaging window $L$ from 1 to 10: The Python code used in this section is selection.py
(19) H(t)h=b(H(t1)h)+U(t).H(t)-h=b(H(t-1)-h)+U(t).
We analyze residuals $U$; results are in Table 6. We conclude that all cases allow us to model $U$ as IID, so (19) is a true autoregression. However, we cannot model $U$ as Gaussian, except for the cases $L=2$ and $L=3$.
We chose $L=10$ to make our research match Shiller CAPE ratio. We think that low $L$ would make our valuation measure fluctuate wildly in the short run, together with annual earnings. We do not need to simulate the residuals $U$ from (19) in our simulations, anyway.
$L$ $R^{2}$ SW JB Skew Kurt L1O L1A
1 17% 2.8% 0.02% 0.55 1.7 0.062 0.604
2 12% 3.1% 25% -0.4 0.18 0.39 0.43
3 11% 20% 12% -0.38 0.71 0.41 0.378
4 10% 5.9% 1.7% -0.31 1.27 0.392 0.546
5 10% 9.1% 3.1% -0.52 0.78 0.446 0.946
6 10% 11% 5.2% -0.48 0.74 0.438 0.753
7 10% 13% 3.5% -0.56 0.65 0.436 0.725
8 9% 10% 7.5% -0.53 0.39 0.302 0.761
9 9% 2.6% 0.8% -0.67 0.73 0.277 0.632
10 10% 0.3% 0.04% -0.82 1.1 0.291 0.608
Table 6. Residuals
In addition, we might include volatility in (19): H(t)h=b(H(t1)h)+V(t)U(t)H(t)-h=b(H(t-1)-h)+V(t)U(t). However, analysis of the ACF for the residuals $U(t)$ shows it works only for short averaging window. To keep it simple, we will focus in this article on the model (19), delegating the other equation to further research.
6. The Main Model
6.1. Overview
We now combine three parts in the Python code main.py
  • the simplified model (10) of three asset classes from Section 4, with only volatility $V$ and rate $R$ as factors
  • the combined model of S&P earnings and returns (18) from Section 5, without international stocks and corporate bonds
  • the term spread $S$ as factor for returns and earnings growth
6.2. Term spread
Since the long-short term spread $S(t)$ can be both positive and negative, we apply the autoregression of order 1 without taking the logarithm:
(20) $S(t)=a_{S}+b_{S}S(t-1)+Z_{S}(t).$
We estimate $a_{S}=0.6437$ and $b_{S}=1-0.4605$. The $p$-values for Student $T$-test for $a_{S}=0$ or for $b_{S}=1$ are less than $0.1\%$.
6.3. Earnings growth
We model earnings growth $G(t)$ as linear regression with respect to change in rate $R$, and term spread $S$:
(21) $G(t)=c_{G}-d_{G}(R(t)-R(t-1))-b_{G}S(t-1)+Z_{G}(t).$
But remember that we divide $G(t)$ by $V(t)$ to make these growth terms closer to Gaussian. Similarly to the switch from (7) to (9), we need to change (21) to
(22) $G(t)=a_{G}+b_{G}S(t-1)+c_{G}V(t)-d_{G}(R(t)-R(t-1))+V(t)Z_{G}(t).$
variable coefficient estimate stderr t P>|>|t|| [0.025, 0.975]
constant $a_{G}$ 0.07756 0.048 1.630 0.107 [-0.017, 0.172]
term spread $b_{G}$ 0.04786 0.018 2.667 0.009 [0.012, 0.083]
volatility $c_{G}$ 0.007841 0.006 -1.370 0.174 [-0.019, 0.004]
rate change -$d_{G}$ 0.03721 0.025 1.466 0.146 [-0.013, 0.088]
Table 7. Earnings growth regression results for (22). Here $R^{2}=13.8\%$. Note $c_{G}$ and $d_{G}$ are not significant, we could remove them. But we chose to retain them, to stress dependence upon the BAA rate and term spread.
Also, we model S&P returns $Q(t)$ as a linear regression versus change in rate R(t)R(t1)R(t)-R(t-1), term spread S(t1)S(t-1), and the valuation measure H(t1)H(t-1). It is important to divide by the annual volatility, as in (9) or (22):
(23) $Q(t)=a_{Q}-k_{Q}H(t-1)-d_{Q}(R(t)-R(t-1))-b_{Q}S(t-1)+c_{Q}V(t)+V(t)Z_{Q}(t).$
variable coefficient estimate stderr t P>|>|t|| [0.025, 0.975]
constant $a_{Q}$ 0.26851 0.032 8.440 0.000 [0.205, 0.332]
term spread -$b_{Q}$ -0.03412 0.012 -2.913 0.004 [-0.057, -0.011]
volatility $c_{Q}$ -0.013568 0.004 -3.523 0.001 [-0.021, -0.006]
rate change -$d_{Q}$ -0.078238 0.017 -4.738 0.000 [-0.111, -0.045]
valuation measure -$k_{Q}$ -0.1644 0.045 -3.667 0.000 [-0.253, -0.075]
Table 8. Domestic returns regression results (23); $R^{2}=53.1\%$. All coefficients are significantly different from zero. Residuals are IID Gaussian.
Same is done for international returns: Although these are non-US stocks, we study their dependence upon US financial factors: The valuation measure $H$, the change in BAA rate, and the term spread.
(24) $I(t)=a_{I}-k_{I}H(t-1)-d_{I}(R(t)-R(t-1))-b_{I}S(t-1)+c_{I}V(t)+V(t)Z_{I}(t).$
Dependence is much weaker, but sometimes still significant. See results in Table 9.
variable coefficient estimate stderr t P>|>|t|| [0.025, 0.975]
constant $a_{I}$ 0.2627 0.077 3.425 0.001 [0.109, 0.417]
term spread -$b_{I}$ 0.0030 0.019 0.160 0.874 [-0.035, 0.041]
volatility $c_{I}$ -0.0186 0.008 -2.336 0.024 [-0.035, -0.003]
rate change -$d_{I}$ -0.0494 0.025 -2.006 0.050 [-0.099, 0.000055]
valuation measure -$k_{I}$ -0.0955 0.083 -1.155 0.254 [-0.262, 0.071]
Table 9. International returns regression results (24); $R^{2}=41.5\%$. Coefficients $b_{I}$ and $k_{I}$ are not significantly different from zero.
We consider a version of (24) without the spread variable: $b_{I}=0$. We see that $R^{2}$ is not improved. We use this simplified version for the financial simulator, see Table 10.
[…]
constant $a_{I}$ 0.2689 0.065 4.108 0.000 [0.137, 0.400]
volatility $c_{I}$ -0.0188 0.008 -2.412 0.019 [-0.034, -0.003]
rate change -$d_{I}$ -0.0514 0.021 -2.447 0.018 [-0.094, -0.009]
valuation measure -$k_{I}$ -0.0941 0.081 -1.156 0.253 [-0.258, 0.069]
Table 10. International returns regression results (24) without the spread variable; $R^{2}=41.5\%$, same as for the general version of this regression. Residuals are IID Gaussian.
6.4. Combined equations for the complete model
Now let us combine all regression equations for returns $Q,I,B$ of three asset classes: Domestic stocks (23), international stocks (24), and domestic bonds (4). We get the following system (25):
(25) $\begin{split}Q(t)&=a_{Q}+b_{Q}V(t)-c_{Q}S(t-1)-d_{Q}(R(t)-R(t-1))-k_{Q}H(t-1)+V(t)Z_{Q}(t);\\ I(t)&=a_{I}+b_{I}V(t)-c_{I}S(t-1)-d_{I}(R(t)-R(t-1))-k_{I}H(t-1)+V(t)Z_{I}(t);\\ B(t)&=0.01R(t-1)-a_{B}-d_{B}(R(t)-R(t-1))+Z_{B}(t).\end{split}$
Also, combine evolution equations (1), (5), (20), (22), (18) (except the first one) of the four factors in (26): volatility $V$, BAA rate $R$, term spread $S$, and the valuation measure $H$.
(26) $\begin{split}\ln V(t)&=a_{V}+b_{V}\ln V(t-1)+Z_{V}(t);\\ \ln R(t)&=a_{R}+b_{R}\ln R(t-1)+Z_{R}(t);\\ S(t)&=a_{S}+b_{S}S(t-1)+Z_{S}(t);\\ G(t)&=a_{G}+b_{G}V(t)-c_{G}S(t-1)-d_{G}(R(t)-R(t-1))+V(t)Z_{G}(t);\\ H(t)&=Q(t)-\overline{G}(t)+c+H(t-1).\end{split}$
In (26), we have one nonrandom (deterministic) equation for $H$. Each another equation contains a series of IID residuals. After we fit the autoregression (with trend) for $H$, we do not use the innovations $U$ of that autoregression. This stands in contrast to (25) where all three equations have random residuals. Thus the entire simulation model (25) and (26) contain only 7 and not 8 series of innovations.
Also, in (26) we define $G(t)=\ln E(t)-\ln E(t-1)$ and $\overline{G}(t)=\ln\overline{E}(t)-\ln\overline{E}(t-1)$ are growth rates for annual earnings and 10-year trailing averaged earnings. Note that $\overline{G}(t)$ is a function of $G(t),E(t-1),\ldots,E(t-10)$.
Each residuals series is IID. But some of them are not Gaussian. Thus it is reasonable to model these residuals as IID. By construction, they have mean zero. Below, we present Table 11 of statistical functions for each series of these residuals. We include the skewness, kurtosis (both being zero for the normal distribution), $p$-values for Shapiro-Wilk and Jarque-Bera normality tests, and the sum of absolute values (the $L_{1}$ norm) of the first 5 lags for ACF of original and absolute values of residuals. We see that all residuals for factors (volatiltiy, spread, rate, growth) are IID but not Gaussian. However, residuals for all three returns are IID Gaussian. The main goal is accomplished: We found a time series model where innovations are what they are supposed to be: IID. However, we need to use some other method (not a multivariate Gaussian distribution) to model them.
Reg Size Stdev Skew Kurt SW JB L1O L1A
$Z_{V}$ 96 0.3644 0.590 0.057 0.009 0.061 0.401 0.237
$Z_{S}$ 97 1.0537 0.475 2.150 0.008 0.000 0.418 0.689
$Z_{R}$ 97 0.1357 0.814 1.409 0.007 0.000 0.177 0.360
$Z_{G}$ 97 0.021 0.614 2.903 0.000 0.000 0.474 0.253
$Z_{Q}$ 97 0.0135 0.039 0.157 0.344 0.940 0.413 0.590
$U$ 97 0.1767 -0.816 1.102 0.003 0.000 0.291 0.608
$Z_{I}$ 55 0.0181 0.073 -0.512 0.621 0.723 0.429 0.635
$Z_{B}$ 52 0.0263 0.193 0.238 0.857 0.800 0.878 0.706
Table 11. Analysis of regression residuals. Skew = skewness, Kurt = kurtosis, SW and JB = Shapiro-Wilk and Jarque-Bera $p$-values, L1O and L1A = L1 norm for the first 5 lags of the ACF for $Z$ and $|Z|$, Length = the size of $Z$
$Z$ $Z_{V}$ $Z_{S}$ $Z_{R}$ $Z_{G}$ $U$ $Z_{Q}$ $Z_{I}$ $Z_{B}$
$Z_{V}$ 1.0000 0.0538 0.2978 -0.1800 -0.4234 -0.0516 -0.0753 0.0725
$Z_{S}$ 0.0538 1.0000 -0.1817 -0.0661 -0.0608 -0.1581 -0.0397 -0.0951
$Z_{R}$ 0.2978 -0.1817 1.0000 -0.1628 -0.5177 -0.0979 -0.0265 -0.1470
$Z_{G}$ -0.1800 -0.0661 -0.1628 1.0000 0.1524 0.1204 0.0724 -0.1878
$U$ -0.4234 -0.0608 -0.5177 0.1524 1.0000 0.7038 0.3228 0.2714
$Z_{Q}$ -0.0516 -0.1581 -0.0979 0.1204 0.7038 1.0000 0.2768 0.2817
$Z_{I}$ -0.0753 -0.0397 -0.0265 0.0724 0.3228 0.2768 1.0000 -0.0948
$Z_{B}$ 0.0725 -0.0951 -0.1470 -0.1878 0.2714 0.2817 -0.0948 1.0000
Table 12. Correlation matrix of series of residuals, see corrMatrix.xlsx
6.5. Main stability results
Consider the complete system of factors and asset returns
(27) $(V,R,S,G,\overline{G},H,Q,I,B),$
evolving as (25) and (26). The system (27) is not Markov. Indeed, the averaged earnings growth $\overline{G}$ depends on the last 11 years of earnings, not the last two years. But we can fix the last 10 years of earnings. Under this condition, the system (27) is Markov: Its value at step $t$ depends only on the value at step t1t-1. Of course, this Markov process is not time homogeneous, since the equations use these initial 10 years of earnings anyway.
Theorem 3.
Assume the series $(Z_{Q}(t),Z_{I}(t),Z_{B}(t),Z_{G}(t),Z_{R}(t),Z_{S}(t),Z_{V}(t))$ in $\mathbb{R}^{7}$ are IID with mean zero. Assume $k_{Q}\in(0,2)$ and $\beta_{R},\beta_{S},\beta_{V}\in(-1,1)$. Then the extended system (27) has a stationary version.
6.6. Proof of the main result
If we prove stationarity of other components, then the stationarity of (4) and (24) are immediate. Stationarity for $V$, $R$, $S$ follows from standard properties of autoregressions, stated in Lemma 2 below. Now rewrite the equations (26):
(28) $\begin{split}H(t)&=H(t-1)(1-k_{Q})+F(t),\\ F(t)&:=a_{Q}-c-d_{Q}(R(t)-R(t-1))-b_{Q}S(t-1)+c_{Q}V(t)+V(t)Z_{Q}(t)-\overline{G}(t).\end{split}$
Define the function $\varphi(x):=(\ln x)_{+}:=\max(\ln x,0)$. It is clear that
(29) $\mathbb{E}|X|<\infty\Rightarrow\mathbb{E}\varphi(|X|)<\infty.$
Apply Lemma 2. From (29), for each of 5 series of residuals $Z_{Q}(t),Z_{G}(t),Z_{R}(t),Z_{S}(t),Z_{V}(t)$,
(30) $\mathbb{E}\varphi(|Z_{A}(t)|)<\infty,\quad A=G,Q,R,S,V.$
Following the main results of [5, 4], we need to show two results: $\ln|1-k_{Q}|<0$ (which trivially follows from $0<k_{Q}<2$) and $\mathbb{E}\varphi(|F(t)|)<\infty$. Using Lemma 1, we reduce it to:
(31) $\mathbb{E}\varphi(R(t))<\infty,\quad\mathbb{E}\varphi(R(t-1))<\infty;\quad\mathbb{E}\varphi(V(t))<\infty;$
(32) $\mathbb{E}\varphi(S(t-1))<\infty;$
(33) $\mathbb{E}\varphi(\overline{G}(t))<\infty.$
The statements (31) follow from Lemma 2. The statement (32) also does, using (29). Finally, let us show (33). It is easy to see that for positive numbers $x_{1},\ldots,x_{n},y_{1},\ldots,y_{n}$, we have:
$\min\left(\frac{x_{1}}{y_{1}},\ldots,\frac{x_{n}}{y_{n}}\right)\leq\frac{x_{1}+\ldots+x_{n}}{y_{1}+\ldots+y_{n}}\leq\max\left(\frac{x_{1}}{y_{1}},\ldots,\frac{x_{n}}{y_{n}}\right).$
Applying the logarithm to both sides, we get:
(34) $\min\left(\ln\frac{x_{1}}{y_{1}},\ldots,\ln\frac{x_{n}}{y_{n}}\right)\leq\ln\frac{x_{1}+\ldots+x_{n}}{y_{1}+\ldots+y_{n}}\leq\max\left(\ln\frac{x_{1}}{y_{1}},\ldots,\ln\frac{x_{n}}{y_{n}}\right).$
However, for real numbers $z_{1},\ldots,z_{n}$, we have:
$\sum_{i=1}^{n}|z_{i}|\leq\min\limits_{i=1,\ldots,n}(z_{i})\leq\max\limits_{i=1,\ldots,n}(z_{i})\leq\sum_{i=1}^{n}|z_{i}|.$
Applying this to (34), we have:
(35) $\sum_{i=1}^{n}\left|\ln\frac{x_{i}}{y_{i}}\right|\leq\ln\frac{x_{1}+\ldots+x_{n}}{y_{1}+\ldots+y_{n}}\leq\sum_{i=1}^{n}\left|\ln\frac{x_{i}}{y_{i}}\right|.$
Finally, let $n=L$, and let $x_{k}=E(t-k)$ and $y_{k}=E(t-k-1)$ for $k=1,\ldots,L$ in (35). Using (18), we get almost surely
(36) $|\overline{G}(t)|\leq\frac{1}{L}\sum\limits_{k=0}^{L-1}|G(t-k)|.$
To show (33), we need to apply Lemma 1 and (36), and finally to show
$\mathbb{E}\varphi(|G(t)|)<\infty.$
But this, in turn, follows from Lemma 1 and (29). Thus the proof of  (33), and with it the main result, is complete.
Lemma 1.
If two random variables $X,Y$ (not necessarily independent) have
(37) $\mathbb{E}\varphi(|X|)<\infty,\quad\mathbb{E}\varphi(|Y|)<\infty,$
then the sum and product of these two random variables also satisfy these properties:
$\mathbb{E}\varphi(|X+Y|)<\infty,\quad\mathbb{E}\varphi(|XY|)<\infty.$
Proof.
First, let us show this for $X+Y$. This function $\varphi$ is nondecreasing but not concave:
$\varphi(x)=\begin{cases}\ln x,\,x\geq 1;\\ 0,\,x\in[0,1].\end{cases}$
But it has a concave hull: the smallest concave $\psi:\mathbb{R}_{+}\to\mathbb{R}$ such that $\psi(x)\geq\varphi(x)$ for all $x\geq 0$. We can find it by drawing a tangent line from $(0,0)$ to $x\mapsto\psi(x)=\ln x$ at the point $(a,\ln a)$ for $a>1$. We compare the slope of this line: $\ln a/a=\psi^{\prime}(a)=1/a$, equivalent to $\ln a=1$ and $a=e$. Therefore,
$\varphi(x)\leq\psi(x)=\begin{cases}x/e,\,x\leq e;\\ \ln x,\,x\geq e.\end{cases}$
This function $\psi$ satisfies
(38) $\psi(x+y)\leq\psi(x)+\psi(y),\quad x,y\geq 0.$
Indeed, from concavity of $\psi$, we have: $\psi(x+y)-\psi(x)\leq\psi(y)-\psi(0)=\psi(y)$. Thus for any random variable $X$,
(39) $\mathbb{E}\varphi(|X|)<\infty\Leftrightarrow\mathbb{E}\psi(|X|)<\infty.$
Applying (39) to (37), we get:
(40) $\mathbb{E}\psi(|X|)<\infty,\quad\mathbb{E}\psi(|Y|)<\infty.$
Applying (38) to (40), we get: $\mathbb{E}\psi(|X+Y|)<\infty$. Finally, apply (39) again and complete the proof that (37) implies $\mathbb{E}\varphi(|X+Y|)<\infty$. Next, we prove this for $XY$. Indeed,
(41) $\varphi(xy)\leq\varphi(x)+\varphi(y),\quad x,y\geq 0.$
To prove (41), we need to consider four cases, comparing $x,y,xy$ with $1$:
  • if $x,y\leq 1$, then $xy\leq 1$, so $\varphi(x)=\varphi(y)=\varphi(xy)=0$;
  • if $x,y\geq 1$, then $xy\geq 1$, so $\varphi(x)=\ln x$, $\varphi(y)=\ln y$, and $\varphi(xy)=\ln(xy)=\ln x+\ln y$;
  • if $x\leq 1$ and $y\geq 1$ so that $xy\leq 1$, then $\varphi(x)=0$ and $\varphi(y)=\ln y$, but
    $\varphi(xy)=0\leq 0+\ln y=\varphi(x)+\varphi(y);$
  • if $x\leq 1$ and $y\geq 1$ so that $xy\geq 1$, then $\varphi(x)=0$ and $\varphi(y)=\ln y$, but
    $\varphi(xy)=\ln(xy)=\ln x+\ln y\leq 0+\ln y=\varphi(x)+\varphi(y).$
This completes the proof of (41), and with it the lemma. ∎
Lemma 2.
Take IID $\delta(t)$ for $t=0,1,2,\ldots$ with finite mean. Fix a $c\in(0,1)$. Then there exists a unique stationary version of the autoregression $Y(t)=cY(t-1)+\delta(t)$, with $\mathbb{E}|Y(t)|<\infty$.
Proof.
This is a well-known result, see for example the classic book [3]: Extending $\delta(t)$ for negative $t$, we solve
$Y(t)=\sum_{k=0}^{\infty}c^{k}\delta(t-k)$
and observe that this series converges in $L^{1}$. Indeed, the $L^{1}$ norm $\delta|\!|_{1}:=\mathbb{E}|\delta(s)|$ is finite and the same for all $s$, so
[…]
This completes the proof of Lemma 2. ∎
7. Financial Simulator
7.1. User interface
We made an online financial app asarantsev.pythonanywhere.com implementing annual simulation of returns of these three asset classes: USA stocks, international stocks, and USA corporate bonds. We use four factors: the valuation measure $H$, the BAA rate $R$, the term spread $S$, and the volatility $V$. The back end is written in Python with Flask framework flask_app.py and the front end is written in HTML in main_page.html In an advanced version of this simulator, we allow to change initial values of these factors. The front end for this option is written in HTML in complete_page.html The residuals are written in Excel file innovations.xlsx
We fix the last 10 years of earnings, 2015–2024. We consider only nominal data: We failed to model inflation-adjusted bond returns with meaningful regressions and IID residuals.
We allow time horizon of 1–50 years. Almost none plans for more than that, and visualizing the simulation then becomes difficult. The stock/bond proportion can be adjusted at the start and at the end (for example, 80/20 now and 60/40 at the end of the simulation, which is 30 years from now). In between, it is interpolated linearly. Within the stock portfolio, the proportion of domestic and international stocks is set by the user and is constant.
The number of simulations is 10000. We rank the simulations by final wealth, and pick the 5 paths which correspond to 10%, 30%, 50% (median), 70%, and 90%. We picked them because we consider these to be representative of the entire distribution of possible paths. It is remarkable to see how much variance is in these returns. The 90%-path final wealth can exceed the 10%-path final wealth by factors of 100, if the time horizon is long enough and the stock market proportion in the portfolio is more than 50%. The output is in Figure 6.
Figure 6. Simulator Output. On the left, we have the graphs of 10%, 30%, 50%, 70%, 90%-ranked paths by final wealth. On the right, we have the written problem input, and output summary statistics.
7.2. White noise simulation
We simulate the $\mathbb{R}^{d}$-valued white noise ($d=7$) using multivariate kernel density estimation (KDE) with Gaussian kernel. Recall that we have N=20241927=97N=2024-1927=97 data points $z(1),\ldots,z(N)$. For the bandwidth, we use the Silverman’s rule of thumb, slightly modified: The covariance matrix of this multivariate Gaussian kernel is diagonal, so de facto it is a product of one-dimensional Gaussian kernels. Now, take each component. Let $s$ be its empirical standard deviation, and let $r$ be its interquartile range: the difference between the 75% and the 25% quantiles. Then the bandwidth is
(42) $\sigma=\left(\frac{4}{d+2}\right)^{1/(d+4)}\cdot N^{-1/(d+4)}\cdot\min\left(s,\frac{r}{1.34}\right).$
We compute $\sigma=\sigma_{i}$ from (42) for each $i$th component. We adjusted the mulivariate Silverman’s rule of thumb to make this bandwidth look more like the univariate one. Thus the density looks like this:
$F(\mathbf{x})=\frac{1}{N}\sum\limits_{k=1}^{N}\prod\limits_{i=1}^{d}\varphi(x_{i},z_{i}(k),\sigma_{i}),$
where $\varphi(y,m,s)$ is the probability density of $\mathcal{N}(m,s^{2})$.
However, some series of residuals do not have full 97 data points, see Table 11. We provide the following method of filling the missing data in the Excel file innovations.xlsx implemented in innovations.py and with output written in another Excel file filled.xlsx
Assume $(X_{0},X_{1},\ldots,X_{d})$ is a(d+1)(d+1)-dimensional IID sample, such that $X_{0}$ does not have full data, but all other components do have complete data. Our algorithm is:
  • Fit the linear regression $X_{0}=a_{0}+a_{1}X_{1}+\ldots+a_{d}X_{d}+Y$ for the data points when all components are present. Compute the residuals $Y$.
  • Sample $Y^{\prime}$ from existing residuals $Y$ randomly, uniformly, with replacement.
  • For the missing data $X_{0}$, simulate it as $X_{0}^{\prime}=a_{0}+a_{1}X_{1}+\ldots+a_{d}X_{d}+Y^{\prime}$, where $Y^{\prime}$ is such sampled residual, and $X_{1},\ldots,X_{d}$ are corresponding (existing) values.
First, we do this for $Z_{V}$, with one missing data; then with $Z_{I}$, and then with $Z_{B}$. At each step, the series which were filled previously (that is, which played the role of $X_{0}$ above) are used as well for linear regression (becoming one of the components $X_{1},\ldots,X_{d}$ above).
We are aware that this is a non-standard statistical procedure. However, we could not find in the literature any other way of doing this. Classic machine learning methods such as nearest neighbors reduce variance. We understand the need for further research, and we welcome any suggestions for other existing methods.
The multivariate kernel density estimation works as follows, using residuals written in the Excel file filled.xlsx
  • For each component $i=1,2,\ldots$, generate $N_{i}\sim\mathcal{N}(0,1)$ IID.
  • Pick a random time $\tau=1928,\ldots,2024$ uniformly, independently of $N_{1},N_{2},\ldots$
  • Generate $Y=Z(\tau)+\begin{bmatrix}\sigma_{1}N_{1}&\sigma_{2}N_{2}&\ldots\end{bmatrix}$
7.3. Withdrawals and contributions
Retirees or parents of college students regularly withdraw some funds from their wealth to support themselves or their children. Savers for future retirement or college regularly contribute to their (often tax-exempt) savings accounts. Endowments and souvereign wealth funds both contribute and withdraw regularly. It is easiest to model annual contributions or withdrawals, since the time step in our simulation is equal to 1 year. If the wealth decreases to 0 or below, we fix this wealth at zero for the rest of the path. We compute the average ruin year for such paths.
We compute the ruin probability: Which proportion of paths end up with zero wealth. Also, for each simulation, we compute which year it went bankrupt, if any. This allows us to display these ruin moments for the chosen five simulations with 10%, 30%, 50%, 70%, 90% ranked final wealth; and the average ruin time for all paths that end in ruin.
However, mostly contributions or withdrawals happen monthly. We allowed this without creating a separate model for simulating monthly (instead of annual) returns. Assume annual returns are simulated already. Then in each simulation, and in each time step of 1 year, we linearly interpolate the wealth process within this year. We use it to compute the resulting change in wealth for monthly withdrawals or contributions. In addition, we allow for the quarterly option (although this is less common).
Figure 7. Advanced simulator output. Initial wealth is 2.5 million, time horizon is 30 years, with initial factors: Volatility 10, BAA rate 3%, Valuation measure -0.5, long-short spread 1.5%. The stock/bond split is 60/40 at the start and 40/60 at the end, and the domestic/international stock split 70/30. Monthly withdrawals, annual initial amount 100K, with 3% increase. On the left, we have the graphs of 10%, 30%, 50%, 70%, 90%-ranked paths by final wealth. On the right, problem input, and output summary statistics.
7.4. Frequent contributions/withdrawals
Assume we allow to contribute or withdraw funds from our portfolio more frequently than annually. The most common cases in practice are monthly or quarterly. We did not invent a complete model monthly or quarterly returns.
Instead, we build on the annual model. Assume we already simulated one path, one annual step of our portfolio with arithmetic returns $r$. Assume that we do not change portfolio weights (stock/bond and domestic/international stocks) during this one year. To simulate frequent regular withdrawals, we make a very primitive assumption: Growth during this year has constant exponential rate. Take $T$ small time steps per one year: $T=4$ for quarters, and $T=12$ for months. Then wealth $1$ at the beginning of this year turns into wealth $(1+r)^{t/T}$ in $t$ small time steps. Annual contributions/withdrawals of wealth $w$, done regularly at equal time intervals, is split in equal amounts $w/T$.
Combine all this together: Contributions of wealth $1$ at small time step $t$ becomes $(1+r)^{1-t/T}$ at the end of this year. Thus the entire sequence of contributions becomes
$\frac{w}{T}\sum\limits_{t=1}^{T}(1+r)^{1-t/T}=\frac{w}{T}\sum\limits_{s=0}^{T-1}(1+r)^{s/T}=\frac{wr}{T}\left((1+r)^{1/T}-1\right)^{-1}.$
We see that the stock/bond allocation makes a difference for the 20-year horizon but not for the 40-year horizon, judging by the ruin probability $P$. In all cases, however, increased bond allocation decreases the variance of a portfolio. In our view, more than 30% ruin probability is unacceptable for retirees. Since a 40-year time is a reasonable horizon for retirees, one needs to rethink or modify the 4% withdrawal rate. We encourage the readers to try our simulator online and simulate various scenarios.
Since we do not allow inflation adjustment (real quantities) and work only with nominal quantities, we needed to compensate for this: A reasonable annual income now would be laughably low in 50 years. We allowed annual increase or decrease (with constant rate) of contributions or withdrawals. We think that setting 4% annual increase is a reasonable compensation for inflation. Indeed, historical annual inflation rate was around 3%, even with the recent burst of inflation at the start of the 2020 decade.
Stock/bond Years $P$ $T$ Average Wealth 90% Wealth
90/10 20 10.15% 15.54 4.8 11.5
90/10 40 30.66% 24.76 48.7 122.6
60/40 20 4.35% 17.12 2.8 5.8
60/40 40 28.2% 28.46 14.4 37.8
40/60 20 1.51% 18.16 2.0 3.6
40/60 40 32.3% 31.1 5.2 14.0
Table 13. Ruin probability $P$, ruin average time $T$, average final wealth, and final wealth corresponding to the 90%-path.
7.5. Retirement rules
In our previous article [13], we considered validity of the classic 4% withdrawal rate: A retired person withdraws 4% of the initial capital the first year, and then keeps withdrawal annual amount constant after adjusting for inflation. We found this rule works well. However, there we could not simulate the earnings growth directly. As discussed in the Introduction, it is an essential part of such wealth simulation. Thus, in [13] we needed to make certain assumptions on earnings growth, which we found reasonable.
In this article, we have a complete and closed model. We can simply simulate this 4% rule straightaway. As discussed above, we need to compensate for inflation by setting annual increases of withdrawals. We set 4% annual increase in withdrawals. We assume the constant stock/bond portfolio split, and the 50/50 split within stock portfolio between domestic and international stocks. We start with the current (as of June 2025) market conditions, so use the main (not advanced) version of the simulator.
We see that the stock/bond allocation makes a difference for the 20-year horizon but not for the 40-year horizon, judging by the ruin probability $P$. In all cases, however, increased bond allocation decreases the variance of a portfolio. In our view, more than 30% ruin probability is unacceptable for retirees. Since a 40-year time is a reasonable horizon for retirees, one needs to rethink or modify the 4% withdrawal rate. We encourage the readers to try our simulator online and simulate various scenarios.
8. Conclusions and Further Research
In this article, we fit several models (from simple to complicated) for wealth invested in portfolios of three asset classes: Domestic stocks, international stocks, and corporate bonds. Our goal for the model selection is:
  • All regression residuals must be IID, and preferably Gaussian;
  • All regression coefficients must be significantly different from zero, or at least not have a large $p$-value for the Student $T$-test;
  • Regressions must have real-world financial sense.
We believe that all models considered in this article satisfy these requirements. We also provide $R^{2}$ values, although we do not consider this goodness-of-fit measure very important. Especially for stock returns, predictive power of various factors measured in $R^{2}$ is usually quite low. Stock returns are close to random walk. We cannot reasonably expect stock indices to have highly predictable returns, otherwise these would be arbitraged away.
However, our research is quite limited. There exist many other more narrow asset classes. For the stocks, there are value/growth factors, small/middle/large size factors, sector/industry portfolios (IT, medical, utilities, railroads), countries (Canada, Germany, China). We can also combine these factors: For example, German small value stocks. For background, see textbooks [9, Chapter 6] or [1, Chapter 7, Section 3].
For the bonds, there are Treasury/corporate/municipal bonds, bonds of various maturities, corporate bonds with given ratings (say CCC), bonds from other countries, and their combinations: For example, junk foreign government bonds of other countries. See [10, Chapter 10, Section 2] or [1, Chapter 9].
We tried to extend our research for some of these asset classes, but failed:
  • Not enough data, sometimes going back only to 1990s and 2000s;
  • Regression residuals are not IID, judging by the ACF.
Further research could try to find other longer data sets extending further to the past, or try nonlinear time series models to ensure residuals are IID.
We also need to include the real (inflation-adjusted) case: As of now, we can do this for stocks but not bonds. Since any financial simulator must absolutely include bonds, we restricted ourselves to the nominal case. See also [1, Chapter 11].
What is more, further research might insert volatility into the bond market analysis, similarly to the stock market returns. In fact, for monthly bond returns, we found that including monthly average VIX improves residuals for bond regressions (makes them closer to IID and Gaussian). For the annual bond returns, this does not seem to work in our case. But we could try other bond classes, maybe for them VIX would work better.
Last but not least, one must improve modeling of monthly/quarterly steps: Change this very primitive model described in Section 7. It seems hard to model monthly returns and monthly changes in factors, at least if you do not consider nonlinear time series. An alternative scenario seems to fix one-year step in one simulation path, and then build a model of monthly contributions/withdrawals around this step. We might need to simulate monthly returns for asset classes which we know, for limited available monthly data.